OpenStudy (anonymous):

Please solve: 1. 35+36 2. (26)3 3. (6*54)÷(2*52)

OpenStudy (anonymous):

Korie? can u do the first one?

OpenStudy (anonymous):

do u know how to add?

OpenStudy (anonymous):

35 +36

OpenStudy (anonymous):

Add the numbers in the second column first 5+6=11

OpenStudy (anonymous):

put the last number from that answer below the 5+6, carry the leftover to the first column

OpenStudy (anonymous):

1 35 +36 ----- 1

OpenStudy (anonymous):

Now add the first column- 3+3+1=7 Put the 7 below the first column. 35 +36 ----- 71 so 35+36=71

OpenStudy (anonymous):

But if you're just learning addition, I think that the others are going to be way too hard for you.